### v2

Abstract
The Intelligent Driver Model (IDM), proposed in 2000, has become a foundational tool in traffic flow modeling, renowned for its simplicity, computational efficiency, and ability to capture diverse traffic dynamics. Over the past 25 years, IDM has significantly advanced car-following theory and found extensive application in intelligent transportation systems, including driver assistance systems and autonomous vehicle control. However, IDM’s deterministic framework and simplified assumptions face limitations in addressing real-world complexities such as stochastic variability, driver heterogeneity, and mixed traffic conditions. This paper provides a systematic review and critical reflection on IDM’s theoretical foundations, academic influence, practical applications, and model extensions. While highlighting IDM’s contributions, we emphasize the need to extend the model into a modular and extensible framework. Future directions include integrating stochastic elements, human behavioral insights, and hybrid modeling approaches that combine physics-based structures with data-driven methodologies. By reimagining IDM as a flexible modeling basis, this paper aims to inspire its continued development to meet the demands of intelligent, connected, and increasingly complex traffic systems.
keywords:
[…]
1  Introduction
Understanding and predicting traffic flow dynamics require accurate modeling of individual vehicle behavior, which serves as a foundation for both theoretical explorations and practical engineering applications in transportation systems. Among the various modeling approaches, car following (CF) models play a pivotal role by describing how drivers adjust their vehicle speed based on interactions with surrounding vehicles. These models provide essential tools for analyzing traffic stability, simulating diverse traffic scenarios, and optimizing intelligent transportation systems (ITS). Over the decades, several classical CF models have been proposed, including the GM model (chandlerTrafficDynamicsStudiesCar1958), the GHR model (gazisNonlinearFollowLeaderModelsTraffic1961), Gipps’ model (gippsBehaviouralCarfollowingModelComputer1981), the Wiedemann model (wiedemann1992microscopic), the Optimal Velocity Model (OVM, (bandoDynamicalModelTrafficCongestion1995)), the Full Velocity Difference Model (FVDM, (jiangFullVelocityDifferenceModel2001)), and Newell’s model (newellSimplifiedCarfollowingTheoryLower2002). Each of these models has contributed to advancing our understanding of driver behavior and traffic flow, offering varying levels of realism, computational efficiency, and theoretical rigor.
Figure 1: The red bars indicate the number of papers citing IDM in the Web of Science between 2000 and 2024. The blue bars indicate relevant papers published in the target list of journals (see Appendix) that at least use IDM for simulation or theoretical analysis rather than just citing it.
Among these, the Intelligent Driver Model (IDM), introduced by treiberCongestedTrafficStatesEmpirical2000, has emerged as one of the most widely adopted CF models in traffic flow research and applications (see Fig. 1). IDM is recognized for its simplicity, computational efficiency, and strong physical interpretability. Unlike earlier models, IDM conceptualizes vehicles as intelligent agents, dynamically adjusting their acceleration based on nonlinear interactions with the leading vehicle, while incorporating factors such as desired speed, safe distance, and relative velocity. This design allows IDM to capture a wide range of traffic states, including free flow, congestion, and stop-and-go oscillations, and to model complex phenomena like phase transitions, hysteresis, and multi-stability. Its versatility has made IDM a benchmark in both theoretical traffic flow studies and practical ITS applications.
Since its introduction in 2000, IDM has been extensively utilized in academic research and real-world applications, see Fig. 1. Table 1 indicates that the IDM has garnered 3458 citations, averaging 138.32 citations per year (CPY). It ranks first in both total citations and CPY among not only classic CF models but also more general models of transportation science. It has facilitated advancements in understanding CF behavior, deriving fundamental diagrams, and analyzing traffic flow stability. Moreover, IDM has been integrated into widely used platforms such as SUMO, enabling its application in evaluating vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication technologies, as well as in testing advanced driver-assistance systems and autonomous vehicle control algorithms. However, despite its widespread adoption, IDM’s deterministic framework and simplified assumptions face challenges in capturing real-world traffic complexities, such as stochastic variability, driver heterogeneity, and the interactions between human-driven vehicles (HVs) and connected and autonomous vehicles (CAVs).
Given IDM’s foundational role in traffic flow modeling and the growing demands of modern transportation systems, this paper aims to systematically evaluate the model’s development, impact, and future potential over the past 25 years. Specifically, this study focuses on:
  • Summarizing IDM’s theoretical foundations and its key contributions to advancing CF theory, providing the basis for understanding traffic flow dynamics.
  • Critically assessing IDM’s limitations and related extensions to address stochasticity, driver heterogeneity, and hybrid modeling approaches, thereby overcoming its shortcomings.
  • Reviewing IDM’s diverse applications, including its contributions to traffic stability analysis, parameter calibration, traffic management, and interdisciplinary domains such as vehicular communications and infrastructure modeling.
  • Proposing future research directions to enhance IDM’s adaptability and relevance in modern traffic systems, focusing on the integration of big data, hybrid modeling methodologies, digital twin technologies, and human-machine interaction.
This paper is organized as follows: Section 2 introduces IDM’s theoretical foundations and key properties. Section 3 discusses IDM’s limitations and extensions. Section 4 reviews IDM’s practical applications, and Section 5 outlines future research directions. Finally, Section 6 concludes with reflections on IDM’s legacy and future potential.
Table 1: Citations of famous classic models included in the Web of Science Core as of May 2025. wiedemann1992microscopic is not included in the Web of Science, so the citation is collected in google scholar database.
Classic Models Citations Classic Models Title Publication Source Citations Per Year(CPY)
Car Following Models 3458 IDM
(treiberCongestedTrafficStatesEmpirical2000) Congested Traffic States in Empirical Observations and Microscopic Simulations Physical Review E 138.32
2628 Optimal Velocity Model
(bandoDynamicalModelTrafficCongestion1995) Dynamical model of traffic congestion and numerical simulation Physical Review E 87.60
1773 Gipps Model
(gippsBehaviouralCarfollowingModelComputer1981) A behavioural CF model for computer simulation Transportation Research Part B 40.30
1318 Full Velocity Difference Model
(bandoDynamicalModelTrafficCongestion1995) Full velocity difference model for a car following theory Physical Review E 54.92
985 GHR Model
(gazisNonlinearFollowLeaderModelsTraffic1961) Nonlinear follow-the-leader models of traffic flow Operations Research 15.39
933 General Motor Model
(chandlerTrafficDynamicsStudiesCar1958) Traffic dynamics: studies in car following Operations Research 13.93
881 Pipes Model(pipesOperationalAnalysisTrafficDynamics1953) An operational analysis of traffic dynamics Journal of applied physics 12.24
827 Newell Nonlinear Model(newellNonlinearEffectsDynamicsCar1961a) Nonlinear effects in the dynamics of car following Operations Research 12.92
641 Newell Simplified Model
(newellSimplifiedCarfollowingTheoryLower2002) A simplified car following theory: a lower order model Transportation Research Part B 27.87
378 Wiedemann Model
(wiedemann1992microscopic) Microscopic traffic simulation: the simulation system MISSION Project ICARUS (V1052) Final Report 11.45
More General Models of Transportation Science 3423 NaSch
(nagelCellularAutomatonModelFreeway1992a) A cellular automaton model for freeway traffic Journal De Physique 103.76
3319 LWR
(lighthillKinematicWavesIITheory1955) On kinematic waves II. a theory of traffic flow on long crowded roads Proceedings of the Royal Society of London 47.41
2276 Cell Transmission Model
(daganzoCellTransmissionModelDynamic1994) The cell transmission model-a dynamic representation of highway traffic consistent with the hydrodynamic theory Transportation Research Part B 73.42
1591 Vickrey’s bottleneck model
(vickreyCongestionTheoryTransportInvestment1969) Congestion theory and transport investment American Economic Review 28.41
Table 2: Abbreviations and full names of concepts in IDM-related research.
Abbreviations Full names
ACC Adaptive Cruise Control
CACC Cooperative Adaptive Cruise Control
HV Human-driven Vehicle
AV Autonomous Vehicle
CV Connected Vehicle
CAV Connected and Autonomous Vehicle
RL Reinforcement Learning
V2X Vehicle to Everything
$F$ Free Flow
CF Car Following
MCMC Markov Chain Monte Carlo
VANET Vehicular Ad-hoc Network
VSL Variable Speed Limits
2  Fundamental Introduction and Analysis of the IDM
This section aims to provide a comprehensive overview of the IDM’s foundational features, focusing on its mathematical structure, behavioral assumptions, and its role as a key model for understanding individual vehicle dynamics and emergent traffic phenomena. We begin by conducting a formal analysis of IDM’s mathematical framework to clarify its core principles and assumptions. Following this, we investigate how IDM captures traffic fluctuations and vehicle interactions, leading to instabilities and emergent flow patterns. Additionally, we analyze spatiotemporal traffic phases, emphasizing IDM’s ability to model complex traffic phenomena across different temporal and spatial scales. For clarity, Table 2 summarizes the abbreviations frequently used in IDM-related research, along with their full names.
2.1 Model Structure and Philosophies
2.1.1 Model Structure
The formulation of basic IDM is as follows:
$a_{n}(t)=a\left[1-\left(\frac{v_{n}(t)}{v_{0}}\right)^{\delta}-\left(\frac{s^{*}_{n}(t)}{s_{n}(t)}\right)^{2}\right]$ (1)
where $a_{n}(t)$ represents the vehicle’s acceleration, $a$ is the maximum acceleration, $v_{n}(t)$ is the current velocity, $v_{0}$ is the desired velocity, $s_{n}(t)$ is the actual spacing, and $s^{*}_{n}(t)$ is the desired dynamic spacing. The exponent $\delta$ determines how quickly acceleration decreases as a vehicle approaches its desired speed. The term $\left(\frac{s^{*}}{s}\right)^{2}$ is used to describe the nonlinear sensitivity of drivers to changes in vehicle spacing, particularly as the actual spacing $s$ approaches the desired spacing $s^{*}$. The quadratic form of this term is closely aligned with Stevens’ power law, which states that the perceived intensity of a stimulus follows a power-law relationship with the actual stimulus. Specifically, the exponent $n=2$ in IDM reflects a balance between capturing drivers’ heightened sensitivity to small spacings and maintaining numerical stability in the model. While a linear form ($n=1$) would underestimate drivers’ rapid responses in critical situations, and a cubic form ($n=3$) would lead to excessive sensitivity, the quadratic term provides a reasonable approximation of how drivers respond nonlinearly to perceived distance changes, as suggested by Stevens’ law for psychophysical scaling (parkDriversCharacteristicsPerceptionLead2001).
The desired spacing $s^{*}_{n}(t)$ is defined as:
$s^{*}_{n}(t)=s_{0}+\max\left(Tv_{n}(t)+\frac{v_{n}(t)\Delta v_{n}(t)}{2\sqrt{ab}},0\right)$ (2)
where $s_{0}$ is the minimum spacing, $T$ is the desired time headway, $\Delta v_{n}(t)=v_{n}(t)-v_{n-1}(t)$ is the velocity difference to the preceding vehicle n1n-1, and $b$ is the comfortable deceleration. The model captures critical driving behaviors by decreasing acceleration as the vehicle approaches its desired speed and ensuring safety through dynamic spacing.
The third term in the desired gap formula, $\frac{v\Delta v}{2\sqrt{ab}}$, reflects the principle of kinematic deceleration. This term is derived from the classical braking distance formula:
$d=\frac{v^{2}}{2a},$ (3)
which quantifies the distance required to stop under constant deceleration. In the IDM, this component ensures that the vehicle maintains a sufficient safety margin when approaching a slower lead vehicle. By incorporating the relative speed $\Delta v$ between vehicles, the model enables smooth and anticipative deceleration, preventing abrupt braking and reducing the risk of collisions. This "intelligent braking strategy" is a core feature of the IDM and one of the key reasons for its ease of use and popularity.
The safety property of the IDM arises directly from the structure of its desired gap function. The model ensures that a following vehicle maintains the desired time headway of $T$, while dynamically reacting to closing speeds through the interaction term. When the actual gap $s$ falls below the desired gap $s^{*}$, the braking term in the acceleration equation becomes dominant, triggering rapid deceleration to prevent potential collisions. This mechanism guarantees collision-free operation as long as the physical braking limits are not exceeded (e.g., under extreme driver behavior or external disturbances). Moreover, $s_{0}+Tv_{n}(t)$ is the classical linear model that gives rise to the triangular fundamental diagram.
The collision-free nature of the IDM has been analytically validated and confirmed through simulation studies (e.g., treiberTrafficFlowDynamics2013). This robust safety property makes IDM particularly well-suited for applications such as Advanced Driver Assistance Systems, where proactive collision avoidance is critical. On the other hand, it makes IDM (or any safe CF model) not suitable for evaluating driving safety.
Beyond its safety-centric design, another key advantage of the IDM lies in its explicit representation of dynamic time headway variations. In the desired spacing formula (2), the term $Tv_{n}(t)$ represents the nominal space, while the additional term $\frac{v_{n}(t)\Delta v_{n}(t)}{2\sqrt{ab}}$ dynamically adjusts the effective spacing based on the closing speed $\Delta v_{n}(t)$. The desired time headway, defined as $s^{*}_{n}(t)/v_{n}(t)$, therefore adapts to both velocity and relative speed variations. This explicit coupling between spacing and velocity not only provides a clear physical interpretation but also enhances the model’s ability to replicate real-world driving behavior under diverse traffic conditions. Such a feature is particularly important for capturing the dynamics of congested or mixed traffic flows.
Moreover, as outlined in Equations (1) and (2), the IDM is a nonlinear CF model characterized by six parameters: $a$, $b$, $v_{0}$, $\delta$, $s_{0}$, and $T$. According to treiberMicroscopicCalibrationValidationCarFollowing2013, these parameters are intuitive, each is related to a specific driving regime: the desired speed $v_{0}$ is relevant for cruising in free-flow scenarios, the desired time gap $T$ pertains to steady-state CF, the minimum gap $s_{0}$ is associated with creeping and standing traffic, and the maximum acceleration $a$ and desired deceleration $b$ relate to non-steady traffic flow. The exponent $\delta$ governs this reduction: a larger $\delta$ delays the reduction in acceleration as the vehicle approaches the desired velocity. In the limit $\delta\to\infty$, the acceleration curve aligns with that of the Gipps model, while $\delta=1$ replicates the overly smooth acceleration behavior characteristic of the optimal velocity model. This design not only gives IDM strong physical meaning but also provides analytical value.
2.1.2 Model Philosophies
In addition to the parameters with obvious physical meanings, IDM also effectively captures the acceleration behavior of drivers by balancing two key components: a non-interaction term and an interaction term. The model defines the desired speed as the maximum attainable speed in free-flow conditions (FF), where drivers can accelerate smoothly to maintain this speed without hindrance. However, in CF scenarios, the driver’s ability to reach the desired speed is constrained by the leading vehicle, although the goal to achieve this speed remains unchanged. The IDM acceleration formula incorporates a non-interaction term, reflecting the driver’s acceleration in the absence of external influences, and an interaction term, which accounts for the deceleration required due to the presence of a leading vehicle. In FF conditions, acceleration is primarily determined by the non-interaction term, while in CF conditions, both terms significantly influence the dynamics. The relative weight of these terms in IDM is dependent on the driver’s current speed, the speed of the leading vehicle, and the distance between them. These factors collectively ensure that the IDM provides a nuanced and realistic depiction of vehicle dynamics across diverse traffic scenarios.
Hence, IDM is widely used not only because it has a clear physical meaning, but also because it was developed with some basic assumptions built on the modeling philosophy (treiberTrafficFlowDynamics2013):
  1. 1.
    Acceleration Conditions: The acceleration must satisfy specific general conditions to ensure a complete model.
  2. 2.
    Equilibrium Distance: The minimum safe following distance to the leading vehicle should be no less than a designated "safe distance," calculated as $s_{0}+vT$, where $s_{0}$ is the minimum gap and $T$ is the time gap to the leading vehicle.
  3. 3.
    Braking Strategy: An intelligent braking approach governs interactions with slower vehicles or obstacles:
    • Under normal conditions, braking is gradual, smoothly reducing deceleration to zero as the vehicle approaches a steady state or stops.
    • In critical situations, the deceleration exceeds comfortable levels until safety is restored, after which it resumes normal deceleration.
  4. 4.
    Smooth Mode Transitions: Transitions between driving modes, such as from acceleration to CF, are smooth. The model ensures that the jerk, or the derivative of acceleration, remains finite, indicating continuous differentiability across variables.
  5. 5.
    Model Parsimony: The model should be as simple as possible, aiding in effective calibration.
2.1.3 Connections with Classic Models
The modeling philosophy of IDM is not an isolated concept but builds upon a rich history of CF theories. Many of its principles, including safety constraints, equilibrium distances, and smooth transitions between driving states, are inspired by or extend ideas from earlier CF models. To fully understand its motivations and distinctive characteristics, it is essential to examine its relationship to foundational models such as the GM model (chandlerTrafficDynamicsStudiesCar1958), the OVM (bandoDynamicalModelTrafficCongestion1995) and the Gipps model (gippsBehaviouralCarfollowingModelComputer1981). IDM synthesizes and extends their strengths while addressing their inherent limitations, resulting in a model that is both comprehensive and robust.
The GM model, one of the earliest CF frameworks, highlights the importance of relative speed (velocity difference) as a key factor influencing driver behavior. Specifically, it assumes that the acceleration of a following vehicle is proportional to the velocity difference and inversely proportional to the headway between vehicles. This foundational approach captures drivers’ tendency to react to both distance and speed differences but oversimplifies real-world behavior by assuming a linear relationship between these variables. Moreover, the GM model does not account for explicit safety constraints, such as minimum headways or emergency braking mechanisms. IDM addresses these shortcomings by introducing a nonlinear interaction term that captures the effects of both relative speed and headway more realistically. Additionally, IDM incorporates explicit safety measures, ensuring collision-free behavior and smooth transitions between free-flow and CF conditions.
The OVM emphasizes the importance of "optimal spacing" in CF behavior. It assumes that drivers adjust their acceleration to minimize the difference between their current speed and a desired speed, which depends solely on the headway to the leading vehicle. This simple and intuitive framework provides insights into equilibrium spacing and CF dynamics. However, the OVM assumes instantaneous driver reactions, neglecting the finite reaction times observed in real-world driving. It also lacks explicit safety constraints, such as minimum headways or braking strategies, making it insufficient for modeling close-following scenarios or sudden deceleration events. IDM improves upon OVM by incorporating safety constraints and modeling the dynamic interactions between relative speed and headway, thereby addressing these deficiencies.
The Gipps model introduces the concept of "safe speed," explicitly ensuring collision-free behavior by predicting the position of the leading vehicle after a reaction time interval and limiting acceleration and deceleration accordingly. This focus on safety is a significant advancement compared to earlier models. However, the Gipps model relies on piecewise-defined functions, which are not smooth or continuous. This lack of smoothness makes it unsuitable for analyzing higher-order dynamics, such as jerk (the rate of change of acceleration), and hinders its ability to simulate smooth transitions between driving states. IDM inherits the focus on safety constraints from the Gipps model but improves upon it with a continuous and differentiable formulation. This smoothness not only ensures finite jerk but also enables realistic transitions between driving modes, such as accelerating to a desired speed or braking to avoid collisions.
The FVDM, which extends the OVM, incorporates the influence of both the velocity difference and the optimal spacing, making it a hybrid model that combines features of the GM and OVM frameworks. While FVDM provides a more realistic description of CF behavior by including velocity differences, its structural design does not inherently ensure accident-free dynamics. This limitation arises because FVDM lacks explicit safety constraints like those in IDM. On the other hand, the simpler structure of FVDM, inherited from the OVM, makes it more amenable to theoretical analysis. IDM, while more complex, ensures accident-free behavior by construction, as it incorporates both safety constraints and a nonlinear interaction term. However, the added complexity of IDM’s structure makes it more challenging to analyze theoretically compared to FVDM.
In summary, IDM bridges the conceptual gaps between these classical models while addressing their specific limitations. From the GM model, IDM inherits the emphasis on velocity differences but refines it with nonlinear dynamics. From the OVM and FVDM, IDM adopts the concept of desired spacing while ensuring safety through explicit constraints. From the Gipps model, it incorporates the idea of safe speeds but improves smoothness and differentiability. This unification of ideas, combined with its ability to model diverse traffic scenarios, explains why IDM has become a cornerstone of CF modeling in the past decades.
2.2 Basic Properties of IDM
Before applying the IDM in traffic simulations and related research, it is critical to understand its fundamental properties to ensure reliable and realistic results. This section focuses on two key aspects: numerical properties (including numerical update schemes and parameter calibration) and traffic stability, emphasizing their relevance for practical applications such as Advanced Driver Assistance Systems and large-scale traffic simulations.
2.2.1 Numerical Schemes of the IDM
For numerical solution schemes of the IDM, treiberComparingNumericalIntegrationSchemes2015 compared four explicit integration methods: Euler method, ballistic update, Heun method (trapezoidal rule), and standard fourth-order Runge-Kutta (RK4). The study found that the ballistic scheme is the best for all practical purposes. Particularly, it is better than Runge-Kutta 4 (RK4) and Euler. The higher consistency order 4 of RK4 is of no use when lane changes or other discrete events are included reducing the overall consistency order to 1. While Euler has the same consistency order 1 as the ballistic scheme, the prefactor of the numerical error is much smaller. Moreover, the ballistic scheme with larger update time gaps could be considered as a special case of the human factor "limited attention span/action points" with an action (i.e., a changed acceleration) only occurring at each timestep. Most researchers and simulators, including SUMO, use this scheme.
From the parameter calibration perspective, treiberMicroscopicCalibrationValidationCarFollowing2013 explored methodologies for calibrating CF models based on IDM. They emphasized criteria for assessing model quality in addition to RMSE, including completeness, robustness, parameter orthogonality, and intuitive parameters aligned with plausible values. The study highlighted that global calibration based on gaps is more reliable than local calibration or those based on speeds or accelerations. Moreover, they found that a sampling rate of 1 Hz suffices, with data completeness and a minimum interval of 300 seconds being crucial for accurate calibration, while data smoothing had minimal impact. sharmaMoreAlwaysBetterImpact2019b found that there is no one-to-one correspondence between IDM parameters and driving regimes. Specifically, the acceleration behavior of the IDM driver in a specific state is controlled by the interaction of multiple IDM parameters.This highlights the importance of considering parameter interactions and dataset characteristics when interpreting IDM calibration results. punzo_we_2015 conducted a comprehensive analysis of IDM parameters, identifying the minimum time headway $T$ as the most influential parameter in determining model performance and stability. Sensitivity analysis revealed that $T$ explains the majority of performance variability, followed by the acceleration exponent $\alpha$ and maximum acceleration $a$. To balance model simplicity and performance, they proposed a simplified IDM configuration by fixing less influential parameters, achieving faster convergence while maintaining comparable accuracy. This aligns with other findings emphasizing the importance of prioritizing key parameters during calibration for robust model performance.
2.2.2 Traffic Stability
From the perspective of traffic flow mechanisms, the IDM has been extensively investigated for its ability to capture traffic instabilities, such as stop-and-go waves, while ensuring safety by avoiding vehicle collisions through its design. These instabilities arise from delays in drivers adjusting their speed to actual conditions, leading to oscillations where vehicles repeatedly decelerate and accelerate rather than maintaining a steady pace. Such oscillations propagate upstream, amplifying congestion, increasing travel times, fuel consumption, and emissions, with significant implications for traffic efficiency and safety.
The ability of the IDM to reproduce these instability mechanisms is rooted in its string stability properties, which depend on parameters such as the sensitivity of the equilibrium speed to gap changes $v^{\prime}_{e}(s)$, the driver’s maximum acceleration $a$, the comfortable deceleration $b$, and the desired time headway $T$. Here $v_{e}(s)$ represents the equilibrium speed that a driver aims to maintain based on the current vehicle gap $s$ under steady-state traffic conditions. Higher sensitivity $v^{\prime}_{e}(s)$ increases instability, while larger values of $a$ and $b$ improve stability by enabling smoother reactions. The desired time headway $T$ is particularly critical, as shorter $T$ values lead to higher instability due to reduced reaction time.
For low speeds $v_{e}\to 0$, the stability condition simplifies to:
$a\geq\frac{s_{0}}{T^{2}},$ (4)
where $s_{0}$ is the minimum gap. This condition indicates that at low speeds, the driver’s maximum acceleration $a$ must exceed a threshold determined by the minimum gap $s_{0}$ and the desired time headway $T$ to maintain traffic flow stability.
Furthermore, the factors influencing string stability can be described as follows:
  • Sensitivity to gap changes $v^{\prime}_{e}(s)$: Higher sensitivity increases the tendency for string instability, as vehicles overreact to changes in the gap.
  • Maximum acceleration $a$ and comfortable deceleration $b$: Larger values enhance stability by allowing drivers to react more smoothly to disturbances.
  • Desired time headway $T$: Shorter time headways lead to increased instability because of reduced reaction times.
The interplay of these parameters determines the overall stability of traffic flow modeled by IDM. For more insights and mathematical formulations, readers are referred to treiberTrafficFlowDynamics2013.
Over the past 25 years, IDM has found wide application in traffic simulations and research precisely because of its ability to model traffic instabilities and stop-and-go waves. Contrary to what is often stated, the stability of the IDM is not a problem for practical applications. In fact, IDM is often parameterized to intentionally reproduce traffic flow instabilities for studying phenomena like congestion and wave propagation. However, if a more stable IDM is required, the acceleration parameter $a$ is the main lever to control stability. The higher the value of $a$, the more stable the IDM (and its derivatives) becomes. For example, realistic values such as $a>1.5\text{m/s}^{2}$ yield a stable IDM under nearly all other plausible parameter combinations. While such configurations will not produce traffic waves, they demonstrate the flexibility of the IDM to adapt to different stability requirements in practice. This adaptability has solidified the IDM’s position as one of the most widely used CF models in traffic simulations.
Despite its practical utility, the IDM does exhibit limitations when compared to empirical observations. For instance, while IDM can reproduce certain instability phenomena, such as upstream-propagating convective instabilities and transitions between convective and absolute instabilities, its predictive accuracy for traffic fluctuations is limited due to inherent oversimplifications. In jiangTrafficExperimentRevealsNature2014, experimental evidence shows that traffic oscillations exhibit a concave growth pattern, where the standard deviation of vehicle acceleration increases concavely as the disturbance propagates upstream in a platoon (see Fig. 2). This finding has been further validated using various datasets, including NGSIM data and controlled experiments, such as the 25- and 51-vehicle platoon experiments in Hefei (jiangExperimentalFeaturesCarfollowingBehavior2015; tianEmpiricalAnalysisSimulationConcave2016; tianCellularAutomatonModelSimulating2016). Together, these studies confirm that the concave growth of traffic oscillations is a universal phenomenon observed under different traffic conditions and experimental settings. In contrast, traditional CF models, such as the IDM, inherently predict a convex growth pattern of oscillations due to the linear instability of their steady-state solutions (wangStabilityAnalysisStochasticLinear2020). This convex growth highlights a critical limitation of IDM, as it fails to capture the empirically observed concave growth, which is essential for accurately modeling traffic flow dynamics.
Moreover, the IDM assumes a deterministic framework that overlooks the stochastic and heterogeneous nature of driver behavior, which plays a significant role in traffic stability. Further insights from tianRoleSpeedAdaptationSpacing2019 suggest that the fundamental mechanism driving traffic fluctuations lies in the competition between speed adaptation and stochastic effects, rather than being solely attributed to reaction delay as assumed in the IDM. Their analysis of CF behavior in a 25-car platoon experiment shows that reaction delay has a negligible effect on fluctuation growth, while the interplay of speed adaptation and stochastic disturbance dominates the dynamics. This evidence challenges IDM’s core assumptions and indicates that its deterministic nature limits its ability to accurately model real-world instability.
Figure 2: Simulation results (solid lines) of the standard deviation $\sigma_{v}$ of the time series of the velocity of each car of IDM from jiangTrafficExperimentRevealsNature2014. The curves represent the simulation results while the markers such as circles represent the experimental results. The different colors represent different experiment runs in which the velocity of leading vehicle $v_{l}$ is 7,15,30,40 and 50 km/h.
2.2.3 Spatiotemporal Patterns Division
Classic traffic theory forms the foundation by classifying traffic into free flow and congested flow, where congested flow encompasses a variety of spatiotemporal traffic patterns. Free flow is characterized by high vehicle speeds and low densities, with minimal interaction between vehicles. Congested flow, on the other hand, includes a range of traffic states such as stop-and-go waves, wide-moving jams, and localized clusters. IDM, through its nonlinear acceleration framework, can reproduce several distinct traffic patterns, including pinned localized clusters (PLC), moving localized clusters (MLC), triggered stop-and-go waves (TSG), oscillatory congested traffic (OCT), and homogeneous congested traffic (HCT) (treiberThreephaseTrafficTheoryTwophase2010). These patterns arise continuously, with smooth transitions between states, as shown in Fig. 3. This flexibility and ability to model traffic instabilities have established IDM as a benchmark model for simulating traffic dynamics within the two-phase framework.
Figure 3: Dynamic phase diagram of on-ramp-induced congested traffic patterns simulated by IDM from treiberThreephaseTrafficTheoryTwophase2010. The abbreviations denote free traffic (FT), pinned localized cluster (PLC), moving localized cluster (MLC), homogeneous congested traffic (HCT), oscillatory congested traffic (OCT), and triggered stop-and-go (TSG) pattern.
3  Critical Discussion of the IDM and Its Extensions
The bibliometric analysis in the previous section provides a comprehensive overview of IDM’s academic impact and identifies key studies that have shaped its development. By synthesizing insights from these core contributions, several critical limitations of IDM have been highlighted. These include its inability to fully capture stochastic variability, driver heterogeneity, and real-time adaptability, which are essential for modeling complex traffic behaviors in real-world scenarios. Addressing these limitations has become a focal point of research, driving the development of diverse extensions to enhance IDM’s functionality and applicability.
Broadly, these extensions can be categorized into three main directions. First, physics-based extensions focus on refining IDM by incorporating additional behavioral or physical principles to better capture human-vehicle interactions. These efforts include both deterministic approaches, which improve the model’s core mechanisms, and stochastic approaches, which introduce randomness to represent driver variability. Second, physics-data-driven hybrid extensions integrate IDM’s interpretability with the flexibility and adaptability of machine learning and statistical techniques, enabling the model to perform effectively in data-rich and complex traffic environments. Lastly, comparative studies with other CF models evaluate IDM’s strengths and weaknesses, offering valuable insights into its positioning within the broader landscape of CF research.
The following sections systematically review these extensions. We begin with physics-based models, discussing both deterministic and stochastic approaches, followed by hybrid methodologies that combine physical and data-driven techniques. Finally, we examine comparative studies, highlighting their role in contextualizing IDM and its extensions in the broader CF research domain.
Table 3: Summary of IDM Extensions, Features, and Use Cases
Model Improvements / Features Use Cases
IDM - Basis for realistic models; calibration reference; testing V2V/V2I communication; benchmark for AI models; physics-informed ML; traffic simulation
IDM+ / IIDM Steady-state time-gap parameter $T$ Same use cases as IDM with a triangular fundamental diagram
IDM-ACC Reduces sensitivity to cut-ins/lane changes; includes jerk control Adaptive cruise control; longitudinal control for autonomous vehicles (validated on test tracks/public roads)
IDMM (IDM with Memory) Considers adaptation to traffic state Models observed traffic flow dynamics
Stochastic IDM Variants (e.g., HDM, 2D-IDM, Action-Point IDM, etc.) Stochasticity, action points, noise, random parameters Reproduce platoon experiments (e.g., concave growth)
Dynamic Time-Gap IDM Variants (e.g., 2D-IDM, Variance-Driven IDM) Risk allostasis, task difficulty homeostasis, indifference region Describe naturalistic data and platooning experiments
Reaction Time IDM Variants (e.g., HDM, Multi-Anticipative IDM) Reaction time, estimation errors, multi-anticipation Realistic modeling of human drivers
IDM-Based Lane-Changing Model (MOBIL) Lane-changing behavior while preserving driver traits Scenarios with lane changes
3.1 Limitations of IDM
Despite its simplicity and computational efficiency, IDM exhibits several critical limitations that hinder its ability to fully capture the complexity of real-world traffic systems. These limitations primarily stem from its deterministic nature, oversimplified behavioral assumptions, and lack of flexibility in addressing diverse traffic scenarios. Key issues include:
  • Omission of Human Factors: IDM does not explicitly incorporate essential human driving characteristics, such as reaction time, estimation errors, and multi-vehicle anticipation. This omission weakens its ability to reproduce realistic driver behavior and traffic stability, particularly under stop-and-go wave conditions.
  • Deterministic Nature: IDM’s deterministic framework fails to reflect the inherent stochastic variability and heterogeneity in driver behavior, limiting its applicability in scenarios characterized by high uncertainty and variability.
  • Inaccurate Oscillation Dynamics: The model predicts convex growth patterns in traffic oscillations, which contradict empirical evidence showing concave growth. This discrepancy reduces its effectiveness in modeling traffic flow instabilities.
  • Steady-State Gap and Deceleration Issues: IDM often produces unrealistic equilibrium gaps and exaggerated deceleration behaviors, particularly at high speeds or during lane changes. This leads to inefficiencies in traffic flow and unrealistic braking patterns.
  • Over-Simplified Assumptions in Complex Scenarios: IDM’s simplified structure limits its ability to model complex scenarios such as merging, lane-changing, and obstacle avoidance, resulting in behaviors that deviate from real-world observations.
  • Limited Adaptability to Modern Traffic Systems: With its lack of mechanisms for cooperative driving or information-sharing, IDM is less effective in modeling connected and autonomous vehicle (CAV) interactions, as well as mixed traffic environments.
These shortcomings underscore the need for advancing IDM through methodical refinements and extensions. Various extended models have been proposed to address these issues by incorporating features such as stochastic variability, human behavioral factors, and enhanced adaptability for modern intelligent transportation systems. Table 3 provides an overview of these IDM extensions, summarizing their addressed limitations and application scenarios.
To further illustrate the advancements in IDM, the following sections will delve into its extensions from two key perspectives: deterministic and stochastic models. Deterministic extensions primarily focus on enhancing the baseline IDM by addressing specific weaknesses, such as reaction time and equilibrium gap issues, while maintaining its computational efficiency. Stochastic extensions, on the other hand, incorporate randomness and human behavioral variability, enabling more realistic modeling in complex and uncertain traffic environments. Together, these extensions represent significant strides in improving IDM’s robustness and applicability across a wide range of traffic scenarios.
3.2 Physics-Based Extensions of IDM
The development of IDM extensions for modeling driver behavior has seen significant advancements, which can be broadly categorized into the deterministic IDM family and stochastic IDM family. Given the rising focus on automated driving, a dedicated subsection is introduced to specifically address IDM extensions tailored for automated driving . Moreover, we also highlight their role in overcoming IDM’s limitations and enhancing applicability in this rapidly growing field.
3.2.1 Deterministic IDM Extensions
Deterministic extensions of IDM focus on addressing specific limitations, such as reaction time, equilibrium gaps, and oscillation dynamics, through modifications grounded in physical principles. As summarized in Table 3, these extensions include models like IDM+, which resolves the misalignment of the time-gap parameter with steady-state conditions, and IDMM (IDM with memory), which incorporates driver adaptation to traffic states.
For example, treiber_delays_2006 emphasized the importance of incorporating reaction time and human factors. While IDM does not directly account for reaction time, it can be extended to include factors such as reaction delays, estimation errors, and multi-vehicle anticipation. Although these factors can undermine traffic stability, they may be compensated by introducing spatial and temporal expectation mechanisms. This explains why IDM and its deterministic extensions, such as the Human Driver Model (HDM), effectively reproduce empirical traffic dynamics despite their simplicity. It further stabilizes traffic flow by increasing oscillation wavelength and reducing the gradient of stop-and-go waves, aligning better with empirical observations. Similarly, kestingAdaptiveCruiseControlDesign2008 extended IDM for simulating adaptive cruise control systems, creating the IDM-ACC, which captures smoother and more stable driving patterns characteristic of ACC-equipped vehicles. Studies such as li_evaluation_2017; li_evaluating_2017 demonstrated IDM-ACC’s ability to evaluate the impact of CACC systems on highway safety and efficiency, showing its robust applicability in automated driving. Despite differences in model components, HDM and IDM-ACC often produce similar patterns due to their shared foundation in IDM’s "intelligent braking strategy" (see subsections 2.1.2 and 2.2.1).
Other deterministic extensions further refine IDM to address specific weaknesses. For instance, Improved IDM (I-IDM) proposed by treiberTrafficFlowDynamics2013 eliminates exaggerated deceleration at speeds exceeding the desired speed $v_{0}$, ensuring equilibrium gaps align with empirical observations. Sigmoid-IDM (chen_sigmoid-based_2024) incorporates a sigmoid function to better handle traffic oscillations, while Task-Difficulty IDM (TD-IDM) (saifuzzamanIncorporatingHumanFactorsCarFollowingModels2014) accounts for task-difficulty considerations to model human factors in heterogeneous environments.
3.2.2 Stochastic IDM Extensions
Stochastic IDM extensions aim to address the deterministic nature of the original IDM by introducing variability to better represent random driver behavior, heterogeneity, and traffic uncertainties. These models enhance IDM’s realism and applicability in scenarios involving traffic instabilities, mixed traffic environments, and connected driving. As summarized in Table 3, stochastic extensions such as 2D-IDM, Action-Point IDM, and other variants effectively capture phenomena like concave oscillation growth and driver heterogeneity.
jiangTrafficExperimentRevealsNature2014 introduced the two-dimensional IDM (2D-IDM), which revealed the concave growth pattern of traffic oscillations through time-varying headway analysis. Building on this, xiong_speed_2022 utilized 2D-IDM to establish a dynamic programming model for recommended speed at isolated signalized intersections in connected environments, characterizing random human driving behavior. Further advancements were made by tianImproved2DIntelligentDriver2016, who proposed the Improved 2D-IDM (I2D-IDM). This model successfully simulated synchronized flow and traffic oscillations by incorporating defensive driving behavior at high speeds, thus improving realism in high-speed scenarios.
To address traffic oscillations, treiberIntelligentDriverModelStochasticity2018 introduced IDM with white noise and IDM with action points, showing that even basic stochastic elements can effectively reproduce observed fluctuation patterns. Similarly, tianRoleSpeedAdaptationSpacing2019 proposed the Stochastic Speed Adaptation Model (SSAM), which utilized the function formulation of desired spacing of IDM (see Equation (2)) and captures critical traffic instability characteristics such as speed adaptation and spacing indifference. By doing so, SSAM successfully reproduces the concave growth patterns observed in empirical data.
Other stochastic extensions have focused on modeling inter-driver and intra-driver variability. For instance, lindorfer_modeling_2018 presented the Enhanced HDM, which incorporates stochastic Wiener processes to model driving errors, variable reaction times, and driver distractions. zhangGenerativeCarFollowingModelConditioned2022c proposed a time-varying IDM that emphasizes both inter-driver and intra-driver heterogeneity, enhancing the model’s ability to simulate real-world variability in human driving behavior.
3.2.3 IDM Extensions for Automated Driving
In addition to the characterization of driving behavior, IDM has also been extended to various aspects of autonomous driving modeling, such as obstacle avoidance, vehicle merging, ecological driving, network attacks, and information error modeling. sharath_enhanced_2020 proposed AV-IDM for two-dimensional motion planning, specifically in mixed traffic scenarios involving obstacle avoidance in which the surrounding obstacles like vehicles and road boundaries are treated as stimuli. chen_safety_2024 developed a model focusing on freeway merging areas, redefining the desired safety distance in IDM to enhance safety during merging under AV environments. ward_probabilistic_2017 proposed a probabilistic extension of IDM for AVs, incorporating normal distribution noise to improve interaction-aware planning in merge scenarios. zhou_impact_2017 introduced the Multi-AV cooperative (MA-C) IDM, integrating cooperative components to assess freeway merging under various AV ratios. zhou_cooperative_2023 introduced the EIDM for cooperative control in mixed traffic, combining multiple vehicle state information to mitigate the negative impacts of HVs. As to eco-driving of AV, dongFlexibleEcoCruisingStrategyConnected2024a proposed a Flexible Eco-Cruising Strategy for CAVs, employing a hierarchical control framework to optimize driving lane planning and speed, thus reducing driving costs in various traffic conditions. When addressing cybersecurity and information errors,wang_modeling_2020 developed the CIDM to model CAV behavior under cyberattacks, utilizing dynamic communication topologies to analyze the impact of manipulated information. berkhahn_traffic_2022 introduced IDMrm, incorporating random misperception (rm) into IDM to model traffic dynamics at intersections with potential perception errors. li_unraveling_2024presented the Dynamic Cooperative IDM (DC-IDM), addressing information errors and their impact on CAVs. liu_connected_2022 proposed a microscopic model for mixed traffic flow unsignalized intersections for autonomous driving, which incorporates the Ornstein–Uhlenbeck process into IDM to model the random perception errors of vehicles at intersections. luo_modeling_2024 proposed the VC-IDM, integrating a generalized force model to analyze platoon behavior and defense mechanisms under cyberattack scenarios.
Summary: The IDM and its extensions owe much of their success to the foundational "intelligent braking strategy," which ensures collision-free behavior when physically possible treiberTrafficFlowDynamics2013. This feature not only enhances IDM’s safety and robustness but also serves as the unifying principle across all deterministic and stochastic variants, making IDM a reliable and flexible framework for both research and practical applications. In deterministic extensions, models like IDM-ACC and HDM refine IDM’s ability to reproduce empirical traffic dynamics, improve stability, and adapt to automated driving scenarios. Meanwhile, stochastic extensions, such as 2D-IDM and SSAM, introduce variability to better capture real-world phenomena like driver heterogeneity, traffic oscillations, and random behavior, thereby enhancing IDM’s realism in diverse traffic conditions.
Despite its strengths, IDM’s deterministic nature limits its ability to fully replicate the complexity of human driving behavior, particularly the inherent randomness and adaptability of drivers in dynamic traffic environments. This limitation becomes especially pronounced in scenarios involving traffic instabilities or stop-and-go waves, where stochastic elements play a critical role in capturing variability and unpredictability.
In conclusion, IDM remains a robust and widely applicable CF model, particularly for automated driving systems where safety and deterministic behavior are paramount. However, to further expand its applicability to human-driven vehicles (HVs), future research should prioritize the integration of stochastic elements and context-specific behavioral adaptations, enabling IDM to better address the variability and complexity of real-world traffic dynamics.
3.3 Hybrid Physics and Data-Driven Extensions
Hybrid physical and data-driven extensions aim to overcome the limitations of both classical and data-driven models. Classical models, such as IDM, accurately represent traffic mechanisms but rely on strong assumptions and struggle with data uncertainties. On the other hand, data-driven methods, while flexible and adaptive, depend heavily on high-quality data and often lack interpretability. By combining these approaches, hybrid models offer a robust framework that balances adaptability with clarity. Typically, the main approach to extend IDM in data-driven vehicle-following modeling is to combine IDM with reinforcement learning (RL) methods and deep learning or machine learning techniques.
3.3.1 Reinforcement Learning-Based Models
Reinforcement learning (RL)-based extensions of IDM aim to improve decision-making and vehicle control by integrating data-driven adaptability with IDM’s physically interpretable structure. For instance, brito_learning_2022 extended IDM using deep reinforcement learning to predict and identify leading vehicles, estimate motion plans, compute control commands (e.g., acceleration), and optimize trajectories in dense traffic, thereby enhancing safety and feasibility. Similarly, caoTrustworthySafetyImprovementAutonomous2022a developed a decision framework for RL, incorporating IDM as a lower performance bound, where autonomous vehicles (AVs) only adopt RL policies if they outperform IDM.
Additionally, baiLongitudinalControlAutomatedVehicles2024a combined Twin Delayed Deep Deterministic Policy Gradient (TD3)-based RL with IDM to achieve improved longitudinal control, safety, and efficiency. kerbel_shared_2024 incorporated IDM’s expected acceleration as a state variable in shared learning to enhance fleet-wide fuel economy and adaptability. Furthermore, hartRobustCarfollowingBasedDeep2024 proposed an RL model with reward functions grounded in IDM’s assumptions, using parameters such as the desired time gap, maximum and comfortable accelerations/decelerations, and desired speed to guide learning.
3.3.2 Machine and Deep Learning-Based Models
The integration of IDM with machine learning (ML) and deep learning (DL) methods has enabled advanced modeling of vehicle behavior by blending data-driven insights with the established physical frameworks of IDM. For example, chen_investigating_2020 developed an LSTD-IDM model that combines long-term and short-term driving characteristics by merging IDM with longitudinal driving data. moPhysicsInformedDeepLearningParadigm2021c proposed a physics-informed deep learning framework (PIDL-CF), blending IDM with deep learning to predict driver acceleration under various conditions.
Hybrid models have also extended IDM with probabilistic and transformer-based approaches. bhattacharyya_hybrid_2022 introduced a hybrid particle-filtering method combined with collaborative IDM extensions. gengPhysicsInformedTransformerModelVehicle2023b presented the Transformer-IDM (PIT-IDM) for enhanced vehicle trajectory prediction by integrating IDM with deep learning techniques. Similarly, kangTrajectoryBasedEmbeddingRandomCoefficients2023b proposed PGM-IDM, combining trajectory-based embeddings with probabilistic graphical models to improve vehicle-following responses.
Further advancements include liuQuantileRegressionPhysicsInformedDeepLearning2023c, which introduced QRPIDL-CF, fusing quantile regression and deep learning with IDM’s physical insights. tang_grounded_2024 proposed a grounded relational reasoning approach that incorporates IDM’s expected distance into the reward function of a Markov decision process, while zhang_bayesian_2024 developed a memory-augmented Bayesian calibration (MA-IDM) to address parameter uncertainties and behavioral differences in IDM, significantly enhancing its predictive accuracy.
Summary: The development of hybrid IDM extensions represents a promising direction in vehicle-following modeling, successfully combining the interpretability of physics-based models with the adaptability of data-driven methods. These approaches improve IDM’s capability to handle complex traffic scenarios, such as dense traffic, heterogeneous driver behavior, and uncertain data inputs. By enhancing IDM with reinforcement learning, hybrid models can dynamically adapt to changing traffic conditions, while deep learning integration enables the modeling of nuanced driver behaviors.
However, hybrid approaches also face several challenges. First, the reliance on high-quality and diverse data for training remains a significant limitation, as the absence of representative data may lead to poor generalization. Second, while combining physical and data-driven models enhances interpretability compared to purely data-driven methods, hybrid models still risk losing physical consistency due to the black-box nature of machine learning components. Finally, the computational cost of training and deploying hybrid models can be prohibitively high, particularly for real-time applications in autonomous driving.
Overall, while hybrid IDM extensions provide a powerful toolkit for addressing the limitations of traditional and data-driven models, future work must focus on balancing interpretability, computational efficiency, and data requirements to ensure their practical applicability.
3.4 Comparisons with Other Models
3.4.1 Deterministic Model Comparisons
Many studies have compared their proposed models with IDM, but relatively few have conducted comprehensive and unified evaluations that fairly compare all models, including IDM. One such study is punzoCalibrationCarFollowingDynamicsAutomated2021, where deterministic models, including IDM, Gipps’ model, and the full velocity difference model (FVDM) with constant-time headway (CTH) and SIGMOID spacing policies, were systematically evaluated across multiple datasets and calibration settings. IDM demonstrated strong and consistent calibration performance, achieving low errors in speed and acceleration while performing well in spacing accuracy. Although Gipps’ model slightly outperformed IDM in reproducing spacing for both human-driven and autonomous vehicle datasets, IDM maintained a more balanced performance across all three measures: spacing, speed, and acceleration. The robustness of IDM across diverse datasets and calibration settings, especially when using Pareto-efficient objective functions such as NRMSE(s,v,a), underscores its versatility as a reliable model for CF dynamics.
3.4.2 Challenges in Comparing with Stochastic Models
While comparisons between deterministic models are well-established, IDM and stochastic models are often compared qualitatively rather than quantitatively. This is because stochastic models inherently produce non-unique outputs, and their calibration processes are complex, involving additional randomness and parameter variability. These factors make direct comparisons with deterministic models like IDM challenging and often unfair. Additionally, the absence of a unified framework for calibrating stochastic models and comparing them with deterministic models hinders systematic evaluation. Such a framework is crucial for enabling fair and consistent comparisons, helping to elucidate the advantages and limitations of stochastic models relative to deterministic counterparts like IDM.
Summary: While IDM has proven to be a versatile and robust model across various datasets and scenarios, its limitations—such as the lack of stochastic variability, over-simplification of human factors, and challenges in adapting to cooperative driving—necessitate further refinement. Comparisons with deterministic models highlight IDM’s balanced performance across key metrics, but its deterministic nature limits its applicability in more variable scenarios. Comparisons with stochastic models remain a significant challenge due to the absence of a standardized evaluation framework. Future research should prioritize developing such frameworks to enable fair and systematic comparisons, thereby advancing the understanding of both deterministic and stochastic CF models.
4  Applications of IDM
This section provides a structured review of the applications of IDM in traffic research and related domains. The review is organized into five thematic areas: traffic stability analysis, calibration frameworks, traffic bottleneck management and control, section and network-level optimization, and interdisciplinary applications. While certain studies may overlap across these themes, each is assigned to the category that best represents its primary focus to ensure clarity and coherence.
The review begins by exploring IDM’s role in traffic stability analysis, focusing on its use in understanding reaction-based instabilities, string stability, and mixed traffic dynamics. Next, the discussion shifts to calibration frameworks, where IDM has been employed to address challenges in parameter estimation, error modeling, and behavioral feature extraction. Applications in bottleneck management, such as intersections and ramp merging, are then highlighted, with IDM widely used to model vehicle interactions under cooperative and signal optimization scenarios. The section further extends to road segment and network-level studies, where IDM is leveraged for evaluating traffic capacity, stability, and emissions under connected and autonomous vehicle (CAV) environments. Finally, the review touches on IDM’s interdisciplinary applications, such as its use in vehicular communication systems and civil engineering for infrastructure load modeling, showcasing its versatility beyond traditional traffic modeling domains.
4.1 Applications in Traffic Stability Analysis
A significant focus has been on understanding traffic stability through IDM. kestingHowReactionTimeUpdate2008 initiated this exploration by analyzing the influence of reaction time and speed adaptation on traffic stability, using IDM to reveal various instability mechanisms. Building on this, jinEquivalenceContinuumCarFollowingModels2016a demonstrated the equivalence between continuum and CF models, illustrating that higher-order IDM can be transformed into continuous models with consistent stability conditions. monteil_linear_2014 advanced the discussion by examining string stability in mixed traffic flows and the impact of multiple anticipations in cooperative CF models. montaninoHomogeneousHeterogeneousTrafficFlows2021b further delved into string stability in mixed and heterogeneous traffic flows, incorporating uncertainties in model parameters. They later proposed a unified modeling framework (montaninoStringStabilityMixedHeterogeneous2021b) to assess stability, considering driver and vehicle heterogeneity and using IDM to verify this framework. bouadiStochasticFactorsStringStability2022b extended these concepts into the realm of stochastic models, investigating string stability in stochastic CF models using IDM with stochastic variants and the generalized Lyapunov equation to establish stability conditions.
Summary: IDM has served as a foundational tool for probing traffic stability, facilitating analytical insights from reaction-based instabilities to string stability in heterogeneous and stochastic contexts. Yet, much of this literature remains confined to linearized scenarios, externally imposed uncertainties, and idealized driver behaviors, limiting its explanatory power in real-world complexity. As traffic systems grow increasingly interconnected and adaptive, future research must re-situate IDM within hybrid, data-informed architectures that move beyond string stability toward capturing emergent, nonlinear, and multi-agent dynamics.
4.2 Applications in Calibration Framework
Research on IDM-based calibration can be categorized into several key areas. Firstly, the development of calibration frameworks focuses on effectively handling Model Performance (MoP), Goodness of Fit (GoF), and trajectory completeness. Secondly, model evaluation involves calibrating various CF models to compare their performance across different metrics. Thirdly, algorithm design aims to create efficient optimization algorithms for calibration processes. Lastly, behavior feature extraction uses IDM-calibrated data to classify different driving behaviors and vehicle types.
For the development of calibration frameworks, building on this focus on error, jin_reducing_2014 explored error accumulation in CF models calibrated with vehicle trajectory data, deriving a dynamic error model based on acceleration. longBiScaleCarFollowingModelCalibration2024a introduced a bi-scale CF model calibration using corridor-level trajectories, highlighting macro measurement error accumulation. To address the validation aspect, treiberValidationTrafficFlowModels2012 developed a quantitative method for calibrating and validating traffic flow models based on the spatiotemporal evolution of congested traffic patterns. In terms of parameter importance and simplification, punzo_we_2015 conducted sensitivity analysis on driving behavior models, finding leader trajectory more important than parameters, and proposed a simplified model with faster convergence. For setting robust calibration objectives, punzoCalibrationCarFollowingDynamicsAutomated2021 proposed robust guidelines for selecting MoP and GoF for calibrating CF dynamics. They used Pareto efficiency and indifference curves for function comparison. Focusing on data integrity, sharmaMoreAlwaysBetterImpact2019b investigated the impact of vehicular trajectory completeness on the calibration and validation of CF models. Finally, addressing behavioral pattern recognition, zhengMultiObjectiveCalibrationFrameworkCapturing2023b proposed a multi-objective calibration framework to capture behavioral patterns of autonomously-driven vehicles, validating with empirical observations of ACC vehicle dynamics. zhouCalibrationStochasticCarFollowing2025a established the calibration framework of stochastic CF models and revealed the essential differences between deterministic and stochastic CF model calibration using various stochastic IDM variants.
As for the development of the optimization algorithms for calibration, rahman_improving_2015 proposed a stochastic calibration method using Bayesian estimation and MCMC simulation. liGlobalOptimizationAlgorithmTrajectory2016b developed a global optimization algorithm combining direct and gradient searches for CF models. zhongCrossEntropyMethodProbabilisticSensitivity2016b introduced a calibration method using the Cross-Entropy Method and sensitivity analysis under noisy data. fardCopulaBasedEstimationDistributionAlgorithm2019b suggested a copula-based algorithm to explore complex parameter interactions in traffic models. wangIntegratedSelfConsistentMacroMicroTraffic2024a created an integrated macro-micro calibration framework using enhanced optimization and deep learning.
In addition to the above theme, calibration is crucial for model evaluation. zhuModelingCarFollowingBehaviorUrban2018b investigated the applicability of CF models for Chinese drivers, highlighting IDM’s effectiveness and portability on Shanghai highways. campi_roundabouts_2024 assessed various models in roundabout scenarios, with IDM excelling in capturing CAV behavior. zhouExperimentalFeaturesEmissionsFuel2023b used four classical emission and fuel consumption models and found IDM effective at the vehicle pair level for energy consumption and emissions predictions. However, at the platoon level, predicted results were almost constant, contrasting with experimental observations.
Based on calibrated IDM, different behavior patterns and vehicle types can be extracted. ossenHeterogeneityCarFollowingBehaviorTheory2011b used data collected by helicopter to calibrate the IDM to identify differences in following behaviors between different driver types, showing significant heterogeneity. makridis_response_2020 estimated the response time of vehicles equipped with ACC by calibrating IDM parameters and found that it was comparable to that of human drivers. hu_autonomous_2023 calibrated IDM parameters to distinguish the empirical following behaviors of AVs and HVs in Waymo data. li_automated_2023 proposed an AV identification algorithm based on IDM calibration results, that is, using error results to distinguish AVs from HVs.
Summary: IDM has played a pivotal role in the development of calibration frameworks, supporting advances in model performance metrics, optimization algorithms, and behavioral feature extraction. From error modeling to cross-scenario validation, IDM has been widely applied due to its analytical tractability and interpretability. However, most calibration efforts still focus on deterministic model variants, while the calibration of stochastic IDM-family models remains notably underdeveloped due to challenges in parameter identifiability, noise disentanglement, and the lack of robust likelihood functions (punzoCalibrationCarFollowingDynamicsAutomated2021). This limits their empirical applicability despite their theoretical promise. Moving forward, calibration should be re-conceptualized not merely as parameter fitting but as a behavioral inference task, requiring methods that can accommodate uncertainty, heterogeneity, and model adaptivity across varying traffic regimes.
4.3 Applications in Traffic Bottleneck Management and Control
Next, we will focus on two typical bottlenecks, intersections and ramps, and summarize the IDM-related research in the field of traffic bottlenecks from the aspects of intersection system evaluation, vehicle trajectory planning and signal optimization, and ramp merging management.
4.3.1 Intersection Systems
Intersection research mainly includes evaluation of the intersection system, vehicle trajectory planning and signal optimization research, as well as the coordinated optimization research of the two. joererVehicularNetworkingPerspectiveEstimating2014a introduced a method to evaluate vehicle collision probability at intersections, using modified traffic simulators based on IDM to assess intelligent transportation system safety, especially when traffic rules are disregarded. schrader_global_2024 employed Sobol global sensitivity analysis to identify how simulation input parameters like IDM’s parameter impact performance metrics of the intersection system. For vehicle trajectory planning related research, IDM is mainly used as a simulation model for background vehicles or manually driven vehicles (chen_human-like_2023; dong_comparative_2022; ghosh_traffic_2022; guoCoordinationConnectedAutomatedVehicles2023a; jiangEcoApproachingIsolatedSignalized2017a; medina_cooperative_2018; stebbinsCharacterisingGreenLightOptimal2017a; wang_learning_2023; xiaoDecisionMakingAutonomousVehiclesRandom2024a; zhangPlatoonCenteredControlEcoDrivingSignalized2023b; zhang_risk-aware_2023).In addition,zhouCooperativeSignalFreeIntersectionControl2022b used IDM as the microscopic simulation model for CAV under the coorperative intersection control framework. For the signal optimization, IDM is also used as the simulation foundation (cao_gain_2022; sohn_waterfilling_2023; yang_enhancing_2024). duCoupledVehicleSignalControlMethod2021b developed an efficient coupled vehicle signal control method for optimizing traffic signal timing and CAV trajectories in mixed traffic environments, using IDM as the artificially driven vehicle model. huang_reservation-based_2023 proposed a reservation-based cooperative ecological driving model for optimizing CAV trajectories and intersection signals at isolated intersections, using background vehicles simulated by IDM. liEcoDrivingSystemElectricVehicles2018a integrated an electric vehicle eco-driving system and signal control in a V2X environment, using IDM to formulate non-connected electric vehicle following rules and a hybrid algorithm for optimization. liang_equitable_2020 combined connected vehicle technology for traffic signal control at isolated intersections, optimizing platoon dissipation based on IDM estimation of vehicle proximity.
4.3.2 Ramp Merging
For ramp scenarios, the main focus is on research related to collaborative merging. baselt_merging_2014; chen_deep_2023; hou_cooperative_2023; ito_coordination_2019; maProgressiveVirtualRiskBasedVehicle2024a; tangNovelHierarchicalCooperativeMerging2022b; xiongManagingMergingCAVLane2022b; zhou_state-constrained_2019 proposed various coorperative merging methods by using IDM to simulate a HV in a mixed traffic merging scenario. guNetworkTrafficInstabilityAutomated2022b use macroscopic or network fundamental diagrams to study the impact of CAVs on network flow destabilization caused by turning and merging maneuvers. They used IDM, HDM and cooperative IDM to simulate the AV, HV and CAV, respectively. silgu_combined_2022 proposed a controller integrating ramp metering and variable speed limits for freeway traffic, tested via SUMO simulations with CACC and HVs. For the system evaluation of ramp, xiaoUnravellingEffectsCooperativeAdaptive2018a analyzed the impact of CACC deactivation on traffic flow at merging bottlenecks, using simulations to evaluate capacity and queue discharge rates.
Summary: IDM has become a foundational simulation tool in bottleneck management, enabling evaluations of intersection safety, signal control strategies, and ramp merging coordination in mixed traffic. Its integration into intersection studies supports both trajectory planning and cooperative signal optimization, while in ramp scenarios, IDM facilitates the modeling of human-driven vehicles in increasingly complex CAV-HV interactions. However, most applications treat IDM as a passive background model rather than an active behavioral hypothesis, often sidestepping the calibration and validation of IDM under high-interaction, high-uncertainty conditions. Particularly in cooperative merging and signal-free intersection control, the use of IDM-family models lacks rigorous behavioral grounding or stochastic robustness. Future work should re-express IDM in these contexts not merely as a simulation agent, but as a control-informed, behaviorally adaptive module, with a stronger emphasis on empirical validation, uncertainty modeling, and multi-agent coordination under real-world constraints.
4.4 Applications in Section and Network Control
Among the research related to the direct application of IDM, the vehicle control especially under the connected and autonomous environment is most popular. We divide these literature into four types: individual longitudinal control (car following), Individual lateral control (lane change), integration of individual lateral and longitudinal coordinated control.
4.4.1 Vehicle Longitudinal and Lateral control
In individual longitudinal control, scholars tend to design the control algorithm of the automatic vehicle based on physical methods or data-driven methods. Therefore, IDM is still used as a vehicle control model for background vehicles, ACC vehicles, manually driven vehicles or some low-connectivity scenarios (juPredictiveCruiseControllerElectric2023a; kamal_efficient_2016; kamalLookAheadDrivingSchemesEfficient2022a; li_deep_2022; li_effect_2015; liu_semantic_2023; liu_potential_2023; liuSpatiotemporalTrajectoryPlanningAutonomous2024a; luLearningDriverSpecificBehaviorOvertaking2018a; ma_energetic_2022; maPersonalizedDrivingBehaviorsFuel2022a; manolis_real_2020; shiConnectedAutomatedVehicleCooperative2021b; shiDeepReinforcementLearningBased2023b; suo_model-free_2024; tangHighwayDecisionMakingMotionPlanning2022a; tas_making_2018; toghi_social_2022; valiente_prediction-aware_2024; wegener_automated_2021; wuFlowModularLearningFramework2022a; wu_hybrid_2023; yan_unified_2023; yangIsolatedIntersectionControlVarious2016a; yuAssessmentDynamicPlatoonInformation2023a; yuanPredictiveEnergyManagementStrategy2019; zhang_stackelberg_2024; zhangNoMoreRoadBullying2023a; zhou_development_2020; milanesModelingCooperativeAutonomousAdaptive2014b; fernandes_platooning_2012; vivekCyberphysicalRisksHackedInternetconnected2019a). Among these, zhangNoMoreRoadBullying2023a and manolis_real_2020 integrated the IDM into the simulation software VISSIM and AIMSUN, respectively.
Compared with longitudinal control, there are relatively few studies on lateral control, and IDM is generally used as a longitudinal control model to study how to perceive, cooperate, make decisions, and adjust in lane change decisions based on the methodologies of game theory, model predictive control, reinforcement learning and etc. like (chen_adversarial_2022; gongGameTheoryBasedDecisionMakingIterative2023a; guIntegratedEcoDrivingAutomationIntelligent2022a; hwang_autonomous_2022; li_bounded_2024; liuMultistepCooperativeLaneChange2024a; liu_interaction-aware_2023; scheel_recurrent_2022; sheng_cooperation-aware_2023; sunberg_improving_2022; wangModelingBoundedRationalityDiscretionary2022a; zhang_game_2020).
For the individual lateral and longitudinal coordinated control, research usually integrate two types of models into one unified framework (kanagaraj_self-driven_2018; li_developing_2023; mullakkal-babu_hybrid_2021).
4.4.2 Evaluation of CAV Systems on Road section
In the road section scale, many research focus on evaluating the performance of the CAV systems. For example,several research evaluate the stability and the capacity of this system (qin_analytical_2021; qinStabilityAnalysisConnectedVehicles2023a; sunStabilityExtensionCarFollowingModel2023b; sunRelationshipCarFollowingString2020b; shang_impacts_2021; talebpourEffectInformationAvailabilityStability2018a; talebpourInfluenceConnectedAutonomousVehicles2016a; wang_stability_2019; yao_linear_2021; zhou_analytical_2021). alhariqiImpactVehicleArrangementMixed2023a; zhou_modeling_2020 evaluated the impact of different following arrangements in a platoon on the road emissions. mullakkal-babu_comparative_2022 evaluated the adverse impact of takeover behavior on mixed traffic flow at different takeover time levels for fully automated vehicles. rahman_longitudinal_2018 explored the impact of CV platooning on highways on longitudinal traffic safety at different market penetration rates, and studied the role of the allocation of CV (i.e., whether they are limited to specific lanes) on traffic safety. wangEgoEfficientLaneChangesConnected2022a studied the impact of lane change on the sytem. wuTimeDependentPerformanceAnalysis80211pBased2020a investigated the impact of disturbances caused by acceleration and deceleration of the lead vehicle, gusts, and uncertainties in the platoon control system, such as aerodynamic drag and rolling resistance torque, on platoon communication. yao_fuel_2021 studied the impact of CAVs on fuel consumption and emissions of mixed highway traffic flows. yaoAnalysisImpactMaximumPlatoon2023a explored the impact of the maximum size of an autonomous platoon on mixed traffic flows. zhou_impact_2021 studied the impact of degradation of CAV platooning control modes on the fundamental diagram. Autonomous driving car testing is an important part of autonomous driving research and development, and IDM also plays an important role in this as the background vehicle (sun_scenario-based_2022; li_scegene_2022; sun_adaptive_2022; cao_autonomous_2023; zhang_accelerated_2024; zhang_bayesian_2024).
4.4.3 Traffic optimization of road section and network
Furthermore, some system management research on the road section and network have been conducted. Some tried to improve the traffic capacity of road section based on platoon control. Most of them use IDM for simulating the HV like ferreira_impact_2012; ge_optimal_2017; heEcoDrivingAdvisoryStrategiesPlatoon2018a; liu_joint_2022; montaninoTrajectoryDataReconstructionSimulationBased2015b; wangRollingHorizonControlFramework2014b; wang_optimal_2022; wang_trajectory_2022, while other platoon control related research always use IDM as the degradation model of CAV. Some use the VSL to optimize the CAV system in which IDM is used to simulate HV or CAV (khondakerVariableSpeedLimitMicroscopic2015a; li_developing_2022; li_integrated_2017; luTD3LVSLLaneLevelVariableSpeed2023a; othman_analysis_2022). mahajan_improving_2024; schakel_improving_2014 proposed a speed-based lane guidance system and a new on-board advisory system, respectively, to provide recommendations on lanes, vehicle speeds, and headway, optimizing lane allocation under high-traffic conditions to reduce capacity degradation. wang_cooperative_2020 studied a cooperative ecological driving system for signal corridors, where the CAV switches between a leading vehicle and a following vehicle, and IDM is used to simulate a CAV with network communication failure.
On the road network scale, various scenarios are simulated with IDM as the underlying model. huang_characterizing_2023 considered different signal control schemes and demand loading patterns, and focused on the overall network level to consider the traffic flow patterns brought about by autonomous driving mixed traffic. claes_decentralized_2011 proposed a distributed vehicle path planning method based on predictive information.
Summary: IDM serves as a foundational component in section and network-level traffic control, frequently used to model human-driven or fallback behaviors in studies of CAV longitudinal-lateral coordination, platoon dynamics, and system-wide optimization. Its integration into mixed traffic simulations underpins evaluations of capacity, stability, safety, and energy consumption across both microscopic and macroscopic scales. However, IDM is often treated as a static behavioral proxy, lacking dynamic adaptation or calibration across diverse control regimes. This simplification becomes especially problematic when IDM is used to simulate degraded CAV behavior or background vehicles in reinforcement learning and game-theoretic frameworks. Moreover, the model’s deterministic structure limits its ability to capture emergent behaviors in stochastic, high-density environments. Future research should reposition IDM not only as a benchmarking tool but as an adaptive module within learning-based, uncertainty-aware, and multi-agent control systems, capable of reflecting the evolving interplay between autonomy, connectivity, and human variability across networked infrastructures.
4.5 Interdisciplinary Applications
IDM is widely recognized for its versatility across numerous research domains beyond its core application in traffic modeling and management of road segments and networks. IDM finds extensive use in areas such as vehicle self-organizing networks, communication algorithms and their associated network attacks within vehicular communications, and even in civil engineering for studies related to bridge loads.
4.5.1 Applications in vehicle communications
In terms of in-vehicle networks, IDM is used as a fundamental model for vehicle operation. wangOptimalFeedbackControlLaw2023a and wang_analytical_2024 have significantly advanced the understanding of cyberattacks on automated vehicles by developing an optimal feedback control law and a comprehensive framework for modeling attacks on adaptive cruise control vehicles. dongDynamicManagerSelectionAssisted2022a introduced a resource allocation algorithm for 5G-V2X platoons, optimizing communication performance under finite block length constraints, crucial for ensuring ultra-reliable low-latency communications. jiaDisturbanceAdaptiveDesignVANETenabledVehicle2014a and jiaNetworkConnectivityPlatoonBasedVehicular2014a delved into VANET-enabled platoon architectures, emphasizing disturbance adaptation and network connectivity, which are vital for effective safety message transmission under varying system parameters. Moreover, segataAutomaticEmergencyBrakingRealistic2013a assessed the integration of automatic emergency braking systems within VANETs, while tianSelfOrganizedRelaySelectionCooperative2017a proposed a decentralized relay selection algorithm based on non-cooperative game theory to enhance cooperative dynamics.
4.5.2 Applications in civil engineering
ge_intelligent_2022 introduced an advanced traffic flow load monitoring system that integrates machine vision and weigh-in-motion systems, leveraging deep learning for precise vehicle detection and extraction of key parameters for traffic flow load simulation. zhou_hybrid_2024 proposed a two-step Hybrid Virtual-Real Traffic Simulation method to replicate the spatiotemporal distribution of bridge loads, demonstrating the broad applicability of IDM in various interdisciplinary fields.
Summary: The adaptability of IDM has enabled its interdisciplinary expansion beyond classical traffic modeling into domains such as vehicular communications, network security, and civil infrastructure assessment. In vehicular networks, IDM supports the evaluation of cyber-physical threats, communication reliability, and cooperative control under V2X and VANET architectures, often serving as the behavioral backbone in simulations of control feedback and message propagation. In civil engineering, IDM-driven traffic simulations are increasingly used to estimate dynamic load effects on infrastructure, bridging traffic theory with structural analysis. Yet, these applications often inherit the behavioral simplifications of the base model, overlooking the contextual nuances of driver behavior under cyber disruptions or infrastructural constraints. To unlock its full interdisciplinary potential, future work should integrate IDM into co-simulation platforms with tighter behavioral-physical coupling, enabling high-fidelity, domain-specific extensions that reflect not only vehicle dynamics but also information flow, structural feedback, and adversarial interactions.
5  Future research directions
Based on the literature review and the analysis of limitations in previous research, several significant gaps and new opportunities for future studies can be identified, as outlined below.
5.1 Combination of Big Data
Current IDM research faces several limitations. Despite the widespread use of open-source datasets like the NGSIM, Naples data set (punzoAnalysisComparisonMicroscopicTraffic2005), Hefei data set (jiangExperimentalFeaturesCarfollowingBehavior2015), and the HighD data set (krajewskiHighDDatasetDroneDataset2018), data scarcity remains a significant challenge, affecting both empirical and experimental data acquisition. The limited availability and scope of these datasets make it difficult to extract universal behavioral or macro-level characteristics. Additionally, insufficient data volume constrains model validation and refinement. This situation not only hampers the development of mechanism-driven models like IDM but also limits the effective training of data-driven models, such as those based on machine learning and deep learning. Issues of data quality and diversity are also prominent, as existing datasets may lack representation of specific traffic conditions or regions. Furthermore, inadequate data collection frequency and precision hinder the ability to capture dynamic traffic changes.
Integrating big data into IDM research can significantly address current limitations through both empirical and experimental data approaches. Empirical data, gathered from traffic monitoring systems, floating car data, and intelligent infrastructure, offers a vast and dynamic dataset. This abundance of data provides a robust foundation for testing and validating models across diverse scenarios. By leveraging these large-scale datasets, IDM can be further refined and integrated with machine learning and deep learning techniques, enhancing traffic prediction capabilities. This integration enables more accurate and adaptive models that can efficiently respond to real-time traffic conditions. On the experimental data front, the advantages are equally compelling. Conducting controlled traffic experiments enables researchers to delve deeper into the mechanisms of driver behavior and traffic flow. Experimental data provides the opportunity to isolate and study specific variables, minimizing interference and uncovering fundamental characteristics and mechanisms (tianCarFollowingDynamicsExperiments2023). This detailed insight is crucial for guiding model development and improving theoretical frameworks. By bridging micro-level decision-making processes with macro-level traffic phenomena, experimental data facilitates a comprehensive understanding of traffic dynamics. This approach not only enriches the theoretical foundation of IDM but also supports the development of models that are both innovative and aligned with real-world complexities.
5.2 Hybrid Modelling Methodology
Hybrid modeling, which combines data-driven and mechanism-driven approaches, offers significant advantages over purely mechanistic or data-driven models. It leverages the strengths of both methodologies: the adaptability and pattern recognition capabilities of data-driven models, and the theoretical rigor and interpretability of mechanism-driven models like (yuanModelingStochasticMicroscopicTraffic2020; yuanMacroscopicTrafficFlowModeling2021). This synergy can lead to more robust and versatile traffic models, capable of addressing complex, real-world scenarios. However, the integration process faces several limitations.
Firstly, the integration often lacks seamless fusion, resulting in limited adaptability and flexibility in practical applications. Secondly, data-driven models and mechanism-driven models often have different underlying structures and assumptions, which makes combining them smoothly a challenge. This can lead to a lack of cohesion in the hybrid model, where components operate in isolation rather than in a unified manner. Additionally, the calibration and tuning processes for each model type can be complex and may not align well, resulting in inefficiencies or conflicts in parameter settings. Additionally, hybrid models can exhibit inconsistent performance across different traffic scenarios, particularly in data-scarce or anomalous situations, affecting their robustness.Moreover, the lack of standardized metrics can lead to unfair comparisons and inconsistent assessments just as wangComparingHundredsMachineLearning2024 pointed out. Furthermore, The challenge of integrating heterogeneous data sources persists, as effectively combining different data types to enhance model performance remains complex. Lastly, computational complexity and real-time applicability are significant constraints; integrated models may demand substantial computational resources, hindering their use in real-time traffic management. Addressing these challenges requires developing more flexible integration frameworks and efficient algorithms to support the advancement and application of IDM models.
5.3 Emerging Technologies
The integration of autonomous driving and V2X technology poses both opportunities and challenges for IDM development. While many studies have explored these technologies, the models developed are often scenario-specific and lack broad applicability. Additionally, there is inconsistency in how vehicle behaviors are defined; IDM is used to simulate HV in some cases, ACC vehicles in others, and even CAV, see Section 3.3. This inconsistency can lead to confusion and limit the effectiveness of simulations. Therefore, there is a pressing need for a universal microscopic behavior model in intelligent connected environments, akin to IDM, that can accurately simulate a variety of vehicle types and support diverse traffic scenarios. Such a model would facilitate more efficient traffic management and optimization in highly interconnected systems.
Digital twin technology offers significant benefits for IDM development by creating high-fidelity digital replicas of real-world traffic systems. This approach allows for the rapid and cost - effective collection of extensive data, which is crucial for model development and training and enhances the accuracy and robustness of IDM. Digital twins enable the simulation of diverse traffic scenarios, providing a platform for rigorous model testing and optimization. Additionally, they support system-wide optimization by identifying and addressing potential traffic bottlenecks and safety issues. Thus, digital twin technology provides essential support for IDM innovation and application, driving advancements in intelligent transportation systems.
5.4 Human-Machine Interaction Integration
The integration of IDM with Human-Machine Interaction (HMI) represents a promising direction for advancing both driver behavior modeling and intelligent transportation systems. As semi-autonomous and autonomous vehicles become more prevalent, the interaction between human drivers and automated systems introduces new challenges and opportunities for IDM development. Traditional IDM focuses on simulating CF behaviors based on deterministic or stochastic rules, yet it often overlooks the dynamic and context-sensitive nature of human decision-making in shared control scenarios. For instance, transitions between manual and automated driving modes, driver reactions to system suggestions, and trust in automation are critical factors that influence vehicle behavior.
Future research should aim to extend IDM by incorporating cognitive and psychological models of driver behavior, enabling more accurate simulation of these interactions. This may involve integrating real-time data from physiological sensors (e.g., eye-tracking, heart rate monitors) and external HMI systems (e.g., dashboards, alerts) to adjust IDM parameters dynamically based on the driver’s state. Furthermore, HMI-integrated IDM can play a pivotal role in designing adaptive interfaces that balance automation and human control, ensuring smooth transitions and reducing driver workload. By capturing the interplay between human decisions and automated systems, such models would not only enhance the realism of traffic simulations but also contribute to safer and more efficient vehicle systems in complex driving environments.
6  Conclusions and Discussion
After a quarter-century of development, IDM has become a cornerstone of microscopic traffic flow modeling. Its conceptual elegance, physical interpretability, and computational efficiency have made it a widely adopted framework in both academic inquiry and practical applications, including traffic simulations, driver assistance systems, and autonomous vehicle control. However, its very ubiquity has also contributed to a certain inertia in the field—dampening critical engagement with its underlying assumptions and limiting the exploration of alternative modeling paradigms. This paper does not aim to offer an uncritical celebration of IDM, but rather seeks to reframe the discourse: to acknowledge its foundational value, while emphasizing the need to re-express and re-engineer the model in light of new technological, behavioral, and methodological demands.
Our analysis reveals that while IDM remains powerful, it is fundamentally incomplete. The model is grounded in a mechanical and rationalist view of driving behavior, which, while effective in controlled scenarios, does not fully capture the complexity of modern traffic systems. As real-world traffic becomes increasingly heterogeneous, stochastic, and influenced by algorithmic decision-making, IDM’s limitations become more apparent. Its deterministic nature, limited treatment of human variability, and unrealistic responses under certain traffic conditions—particularly during congestion or at high speeds—suggest that IDM, in its current form, is misaligned with the realities of contemporary mobility ecosystems.
Thus, we contend that the future of IDM lies not in preserving it as a fixed standard, but in redefining it as a basis for innovation. IDM should no longer be seen as a monolithic solution, but as a modular and extensible framework—one that can be systematically expanded with stochastic elements, integrated with human behavioral insights, and enhanced through hybrid modeling that combines physics-based structure with data-driven learning. Promising directions include the development of context-specific extensions tailored to autonomous vehicles, cooperative driving systems, and eco-driving applications. In this light, IDM’s role should evolve from being a standalone model to becoming an interpretable backbone within a broader ecosystem of intelligent transportation modeling tools—serving not just as a simulation engine, but as a conceptual foundation for next-generation traffic systems.
Furthermore, our review identifies a pressing methodological gap: the absence of standardized, rigorous frameworks for evaluating and comparing CF models across deterministic and stochastic paradigms. Without such benchmarks, IDM’s continued dominance may reflect historical momentum rather than actual superiority. Advancing traffic flow theory requires not only more sophisticated models, but also a clearer understanding of what constitutes model robustness, generalizability, and practical utility under diverse real-world conditions.
Ultimately, this paper does not seek to close the chapter on IDM, but to open a more critical and generative conversation about its future. We hope to shift the perspective from viewing IDM as a fixed modeling artifact to engaging with it as a living theory-in-practice, one that must adapt to meet the demands of intelligent, connected, and increasingly unpredictable traffic environments. IDM’s legacy is not just as a model that captured the past, but as a lens through which we must now reimagine what CF modeling can and should become.
In this sense, IDM is not the end of a modeling journey—it is the beginning of a new phase of theoretical exploration and methodological innovation.
7  Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 72222021, W2411064, 72288101, 71931002), Young Scientific and Technological Talents (Level One) in Tianjin (Grant No. QN20230104) and Beijing-Tianjin-Hebei Basic Research Cooperation Project (Grant No. G2024210009).
Appendix
Table 4 is a list of target journals included in the literature review. We have combined the list of punzoCalibrationCarFollowingDynamicsAutomated2021 and related engineering research community are considered to represent the state of the art in the field.
Table 4: Target Journals Included in the Literature Review.
Category Journal Name
Transportation Science and Engineering Transportation Science
Transportation Research Part A: Policy and Practice
Transportation Research Part B: Methodological
Transportation Research Part C: Emerging Technologies
Transportation Research Part D: Transport and Environment
Transportation Research Part E: Logistics and Transportation Review
Transportation Research Part F: Traffic Psychology and Behaviour
Accident Analysis and Prevention
Transportmetrica A: Transport Science
Transportmetrica B: Transport Dynamics
Electrical and Computer Engineering IEEE Transactions on Intelligent Transportation Systems
IEEE Transactions on Intelligent Vehicles
IEEE Transactions on Vehicular Technology
IEEE Transactions on Automation Science and Engineering
IEEE Transactions on Robotics
IEEE Transactions on Transportation Electrification
Vehicular Communications
Energy and Environment Applied Energy
Energy
Physics and Interdisciplinary Studies Physical Review E
Physica A: Statistical Mechanics and Its Applications
Plos One
[…]